OpenStudy (anonymous):

A standard tennis ball can is a cylinder that holds three tennis balls. If the radius of a tennis ball is 3.5 cm, what percent of the can is occupied by air, not including the air inside the balls?

OpenStudy (anonymous):

the radius of the can is the same as the balls, and the height is 3 ball diameters. find the volume of the cylinder, and subtract the volume of the 3 balls.

myininaya (myininaya):

VC=Volume of cylinder=h*pi*r^2 where h=height and r=radius of cylinder VS=Volume of sphere=4/3*pi*s^3 where s is the radius of the sphere (VC-3*VS)/VC is the percent occupied by air

OpenStudy (anonymous):

Use the formula for volume of a sphere to find the volume of one ball, with radius 3.5mm. Multiply by 3 and you have the total volume occupied by the balls. Then for your sphere volume its gonna be base * height, where the base has a radius of 3.5cm and its height is 6 * r (because there are 3 balls). Subtract total sphere volume from cylinder volume, and you have your answer.

OpenStudy (anonymous):

h = 3(2*3.5)

myininaya (myininaya):

right hick

myininaya (myininaya):

s=3.5 r=3.5

myininaya (myininaya):

snogging just use the formula i gave you and plug in the values hick and i gave you